40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7983     Entry time: Fri Feb 1 12:34:55 2013
Author: Jenne 
Type: Update 
Category: PSL 
Subject: Too much power injected into vacuum 

I noticed (while relocking the MC after Jamie and I zeroed the LSC offsets) that the MC refl power was 4.8 V.  Usually we should be ~4.2, so I closed the PSL shutter and went in to measure the power.  We were injecting ~125mW or a little more.  I had adjusted the power the other day, and through yesterday, it looked fine, but overnight it looks like it drifted up.

ELOG V3.1.3-